OpenStudy (summersnow8):

find all solutions: cos 3x - cos x = 0

OpenStudy (luigi0210):

Factor

OpenStudy (summersnow8):

cos x (3-1)=0?

OpenStudy (luigi0210):

\[\cos^3x-cosx=cosx(\cos^2x-1)\]

OpenStudy (luigi0210):

\[\cos^2x-1=0\]\[cosx=0\]

OpenStudy (summersnow8):

well thats not the answer.... it is suppose to be intervals

OpenStudy (luigi0210):

Yes, can solve it from there?

OpenStudy (luigi0210):

and do you mean from (0, 2pi]?

OpenStudy (summersnow8):

i dont really understand how you factored that....

OpenStudy (summersnow8):

the answer said its x= pi k, pi/2 + pi k

OpenStudy (luigi0210):

It is cos^3x-cosx right?

OpenStudy (summersnow8):

no.... it is simply Cos 3x - Cos x=0

OpenStudy (luigi0210):

>.<

OpenStudy (loser66):

lol,

OpenStudy (luigi0210):

Well I'm not much help then.. 66 can take care of you from here.

OpenStudy (summersnow8):

he left

OpenStudy (loser66):

I don't know whether I can help or not, let my try

OpenStudy (loser66):

I figure out if i go on that way, I will drive you crazy. give me time.

OpenStudy (summersnow8):

hmmm, I don't really understand. thanks